OpenStudy (anonymous):

u = -4i + 1j and v = 4i + 1j; Find ‖u+v‖ ? much help needed. Find the specified vector or scalar.

OpenStudy (anonymous):

\[u+v=-4i+1j+4i+1j=-4i+4i+1j+1j=0i+2j\] from here use the following equation \[||ai+bj||=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

\[\left| u+v \right|=\sqrt{0^2+2^2}=2\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw did u understand?

OpenStudy (anonymous):

sum of two vectors is always a vector.

OpenStudy (anonymous):

that formual comes from pythagorean theorem(for magnitude):